OpenStudy (anonymous):

Geometry - Note: All the letters are angles inside of a pentagon. A,B,C,D are congruent. The ratio of A:E is 2:1 . What is the measure of Angle E?

OpenStudy (anonymous):

E has to be half of the other numbers, and the other numbers are all congruent.

OpenStudy (anonymous):

Scale factor is 2

OpenStudy (anonymous):

Angle E must be 1/2 of all the other angles.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You should have these 5 angles A = x B = x C = x D = x E = x/2 add them up and set that sum equal to 540 and solve for x

OpenStudy (anonymous):

Add them up? What will I be adding up?

OpenStudy (anonymous):

Is it 4x + x/2 = 540? @jim_thompson5910

OpenStudy (anonymous):

Ohh, X = 120 !

jimthompson5910 (jim_thompson5910):

correct, x = 120 so x/2 = 60

jimthompson5910 (jim_thompson5910):

which is the measure of angle E

jimthompson5910 (jim_thompson5910):

I got 540 from this formula S = 180(n-2) n = 5 in this case

OpenStudy (anonymous):

Yep, I got it.